# Adapting LLaMA Decoder to Vision Transformer

### v2

@@ Abstract
This work examines whether decoder-only Transformers such as LLaMA, which were originally designed for large language models (LLMs), can be adapted to the computer vision field. We first “LLaMAfy" a standard ViT step-by-step to align with LLaMA’s architecture, and find that directly applying a causal mask to the self-attention brings an attention collapse issue, resulting in the failure to the network training. We suggest to reposition the class token behind the image tokens with a post-sequence class token technique to overcome this challenge, enabling causal self-attention to efficiently capture the entire image’s information. Additionally, we develop a soft mask strategy that gradually introduces a causal mask to the self-attention at the onset of training to facilitate the optimization behavior. The tailored model, dubbed as image LLaMA (iLLaMA), is akin to LLaMA in architecture and enables direct supervised learning. Its causal self-attention boosts computational efficiency and learns complex representation by elevating attention map ranks. iLLaMA rivals the performance with its encoder-only counterparts, achieving 75.1% ImageNet top-1 accuracy with only 5.7M parameters. Scaling the model to $\sim$310M and pre-training on ImageNet-21K further enhances the accuracy to 86.0%. Extensive experiments demonstrate iLLaMA’s reliable properties: calibration, shape-texture bias, quantization compatibility, ADE20K segmentation and CIFAR transfer learning. We hope our study can kindle fresh views to visual model design in the wave of LLMs. Pre-trained models and codes are available here.
@@ 1 Introduction
The year 2024 saw the meteoric rise of large language models (LLMs), as well as the 4th anniversary of the Vision Transformer (ViT) dosovitskiy2020image . Born in 2020, ViT has revolutionized the computer vision field by adapting the Transformer vaswani2017attention , which was originally designed for natural language, to process images as non-overlapping patches. Notably, ViT design was influenced by the prevailing encoder-only text Transformers at the time, such as BERT devlin2018bert , etc. Accordingly, ViT is allowed to borrow the encoder-only design, i.e., self-attention do not use any causal mask. As a result, advanced vision backbones touvron2021training ; wang2021pyramid ; guo2022cmt and training paradigms bao2021beit ; he2022masked have followed such convention by default.
Meanwhile, the development of text Transformers did not stand still. A series of LLMs with a decoder-only architecture, such as LLaMA touvron2023llama ; touvron2023llama2 , OPT zhang2205opt , and PaLM chowdhery2022palm , have sparked a new wave. Pre-trained decoder-only Transformers have demonstrated remarkable scalability with respect to model capacity and training data in diverse textual tasks. This revolution naturally raises a new issue, i.e., the misalignment between encoder-only visual and decoder-only textual architecture.
@@
In the era of LLMs, could decoder-only Transformers be revitalized in the vision field? In this study, we demonstrate that through straightforward supervised learning, LLaMA architecture itself can process input images with simple yet crucial modifications. We start by modifying a standard encoder-only ViT (e.g., ViT-T/16), progressively adapting its components to align with those in LLaMA. In practice, we observe an attention collapse issue, i.e., the training loss fails to converge by directly adding a causal mask to the attention map. The causal mask restricts the class token from accessing the image’s global information, thereby hindering the optimization of the training loss. To this end, we propose a post-sequence class token technique, repositioning the class token to the end of image tokens (details in Section 3.3). As a result, causal mask can keep the attention score between the class token and others, allowing the model to optimize stably. We also evaluate the advantages of the causal self-attention in reducing computational complexity and enhancing the attention map rank.
@@
Figure 1: Left: iLLaMA architecture. Right: the design roadmap. Colored and gray bars represent the results of the tiny and base regimes, with the red line depicting the training loss of the tiny regime. iLLaMA strives to process visual tokens using standard LLaMa components, e.g., causal self-attention. The proposed PS [cls] and soft mask strategy help overcome training challenges. Block details of ViT dosovitskiy2020image , VisionLLaMA chu2024visionllama , and our iLLaMA is compared in Figure 5 in Appendix A.
@@
Equipped with such modifications, we propose a non-autoregressive decoder-only vision Transformer with causal self-attention inside, dubbed image LLaMA (iLLaMA), as shown in Figure 1. We provide a block level comparison of ViT dosovitskiy2020image , VisionLLaMA chu2024visionllama , and our iLLaMA in Figure 5 in Appendix A. iLLaMA process visual input with our slightly modified causal self-attention. We conduct a comprehensive evaluation of iLLaMA’s properties, including ImageNet-1K classification deng2009imagenet , calibration, shape-texture bias, quantization compatibility, ADE20K semantic segmentation zhou2019semantic , and CIFAR transfer learning krizhevsky2009learning . Experimental results show that iLLaMA delivers favorable and reliable performance to its encoder-only counterparts (i.e., ViT, VisionLLaMA), while maintaining a pure decoder design. More importantly, a spectral analysis on the attention map empirically shows that compared to bi-directional counterparts, causal self-attention has a higher rank (see Figure 4), which allows for learning complex image representation. We hope our work to inspire a re-evaluation of vision backbone design in the era of LLMs and provide fresh insights for their architectural unification.
@@ 2 Preliminaries and Motivation
Autoregressive (AT) and non-autoregressive (NAT) models. AT models have demonstrated remarkable generation ability in LLMs touvron2023llama ; touvron2023llama2 . Besides the text domain, AT models have also been widely used in image van2016pixel ; van2016conditional ; salimans2017pixelcnn++ ; razavi2019generating ; chen2020generative and audio oord2016wavenet tasks. Unlike the stepwise computation paradigm of the AT models, NAT models zhou2019understanding ; saharia2020non ; chan2019kermit ; gu2017non ; gu2019levenshtein generate all output tokens in parallel, thereby controlling the output length easily and reducing inference latency. Our iLLaMA allows for one-step inference and consists entirely of NAT decoders, making it suited for visual perception tasks.
@@ 3 A Roadmap: Solving Attention Collapse and Optimization Improvement
This section introduces the design roadmap of iLLaMA. As we aim to adapt LLMs to vision, we choose LLaMA touvron2023llama and ViT dosovitskiy2020image as language and vision references in light of their successful practices. The trajectory can be divided into two dimensions, i.e., architecture (Section 3.1-3.4) and training techniques (Section 3.5-3.6). First, we focus on block designs including 1) feed foward network, 2) normalization layer, 3) self-attention, 4) positional embedding, illustrated in Figure 1. Next, we study training techniques and develop a soft mask strategy to facilitate optimization. Finally, we provide an analysis in terms of efficiency and attention map rank (Section 3.7). We start with ViT-T/16 and ViT-B/16 with around 5.7M and 86.4M parameters, respectively, and gradually replace the corresponding components with those from LLaMA. We conduct experiments on ImageNet-1K deng2009imagenet , following the training recipe adopted from liu2023dropout (details in Appendix B.1). Considering the differences between visual perception and text generation tasks, we maintain ViT’s non-autoregressive manner in our network. Each step change and the corresponding results are reported in Appendix C.
@@ 3.1 Feed Forward Network (FFN)
FFN structure in Transformer are implemented differently in ViT and LLaMa, i.e., multi-layer perceptron (MLP) and SwiGLU shazeer2020glu . MLP consists of two sequential linear mappings, with a GELU hendrycks2016gaussian function inserted. Meanwhile, SwiGLU combines three linear mappings, integrating a SiLU hendrycks2016gaussian ; elfwing2018sigmoid ; ramachandran2017searching function. This structure allows for the modulation of high-dimensional features through a gating mechanism before reduction. We substituted the Transformer’s MLPs with SwiGLUs, while maintaining comparable computational load. As illustrated in Figure 1, this improves performance from $73.8\%$ to $74.3\%$, and from $81.3\%$ to $82.0\%$ for the ViT-T/16 and ViT-B/16 regime. This improvement highlights SwiGLU’s effectiveness not only in language models but also in vision, inspiring further exploration of other components.
@@ 3.2 Normalization Layer
Transformers need normalization layer for stable training, i.e., layer normalization (LN) ba2016layer in ViT and root mean square layer normalization (RMSNorm) zhang2019root in LLaMA, respectively. We replaced all LNs with RMSNorms in our network and empirically observed that the accuracy of the ViT-T/16 regime increased from $74.3\%$ to $74.5\%$. However, similar improvements in precision were not observed in the ViT-B/16 regime (from $82.0\%$ to $81.7\%$). Nonetheless, compared to LN, RMSNorm removes the shift term computation, bringing simplicity to the network touvron2023llama2 ; vicuna2023 ; roziere2023code ; jiang2023mistral .
@@
   (a) causal mask         (b) causal mask w/ PS [cls]        (c) modified causal mask
Figure 2: Illustration of different masks in self-attention. From left to right: mask in causal self-attention, mask in causal self-attention with the PS [cls] method, and a modified causal mask.
@@
Attention collapse issue. As a common practice for Transformer decoders, the key component for causal self-attention is the causal mask, i.e., a lower triangular mask matrix, illustrated in Eq. 1 and Figure 2(a). With such, each token can get the attention score of all its previous ones. We add the causal mask to our network via a non-autoregressive way. The reason is that visual perception tasks, unlike text generation, require only inference once. As a result, we observe that the training loss fails to converge in both ViT-T/16 and ViT-B/16 regimes (line 1 in Table 1). We posit that such issue stems from the influence of the lower triangular matrix, which prevents the class token from “seeing" other image tokens. As illustrated in Figure 2(a), when the class token is positioned at the start of the patch embedding, its attention score for all other image tokens gets zero due to a causal mask. We term this occurrence as the attention collapse issue, which leads to a loss of connection between the class token and other image patches, thereby hindering the optimization of the network.
@@
To test our hypothesis about the reason of the attention collapse issue, we also explore a mask setting in Figure 2(c). In this setting, we do not change the position of the class token. Instead, we unmask the first row of the mask (i.e., attention score of the class token) on the basis of the causal self-attention, termed as “modified causal mask". Ablation results (line 3 in Table 1) shows that both settings can solve the attention collapse issue as expected, and the “modified causal mask" leads to a better $72.5\%$ accuracy for ViT-T/16 regime, validating our hypothesis about the reason. Although the results do not surpass the performance of bi-directional counterpart, they demonstrate the potential for optimizing causal self-attention in a decoder-only image model. We also observe that the PS [cls] method yields higher accuracy with a slightly larger training loss for ViT-B/16 regime, suggesting lower overfitting.
@@ 3.4 Positional Embedding
A standard ViT use learnable positional embedding (LPE) to preserve positional information, typically adding it directly to the patch embedding. Meanwhile, rotary positional embedding (RoPE) su2024roformer is widely employed in LLMs touvron2023llama ; touvron2023llama2 , which functions within the attention of each block. We first use RoPE alone, which boosts the accuracy of ViT-T/16 and ViT-B/16 regimes to $72.6\%$ and $81.2\%$, respectively. The encouraging results illustrate that the concepts of "position" in image and text do not exist an inherent gap. Since LPE functions only once before all Transformer blocks, keeping it does not disrupt the alignment with LLaMA within each block. Thus, we reintroduce the LPE, which improves the accuracy of ViT-T/16 regime to $73.2\%$, suggesting that the two positional embeddings are not redundant but rather synergistic, contributing to the network performance.
@@ 3.5 Data Augmentation
Mixup zhang2017mixup and cutmix yun2019cutmix that we used to train our iLLaMA (0.8 and 1.0), are borrowed from DeiT touvron2021training ’s recipe. Unlike the bi-directional self-attention used in DeiT, causal self-attention affects the connection between image tokens. Meanwhile, these two hyper-parameters affect the content of the input image, which further influences the subsequent embedding. Thus, we reevaluate their impact on iLLaMA optimization. Specifically, we discover that a combination of 0.1 mixup and 0.1 cutmix improves the performance of the iLLaMA-T/16 to $74.3\%$, whereas a combination of 0.95 and 1.0 leads the iLLaMA-B/16 to a $81.3\%$ accuracy. Other ablations are detailed in Section 4.1.
@@ 3.6 Soft Mask Strategy: Optimization Improvement
Figure 3: Left: soft mask gradually transitions from a bi-directional mask into a causal mask during training through a constant or linear schedule. Right: ablation training loss and test accuracy.
where $i,j\in[1,N]$, ${\bf S}\in\mathbb{R}^{N\times N}$ denotes the soft mask, which is defined as a linear combination of a bi-directional mask ${\bf B}$ and a causal mask ${\bf C}$. $\alpha$ is the hyper-parameter controlling the mask configuration, i.e., soft mask degenerates into ${\bf B}$ or ${\bf C}$ when $\alpha=1$ or $\alpha=0$, respectively. As illustrated in Figure 3(a), $\alpha$ involves three related hyper-parameters: 1) scheme: how $\alpha$ drops from 1 to 0: we try a linear or a constant scheme. 2) cutoff epochs: when will $\alpha$ drops to 0. 3) learning rate (lr) warmup he2016deep ; goyal2017accurate : this hyper-parameter overlaps with the duration of soft mask. We initially set the lr warmup epochs at 50, consistent with previous settings. When using a linear scheme with 50 and 25 cutoff epochs, we observe an improvement in performance for both iLLaMA-T/16 and iLLaMA-B/16 models, achieving $74.9\%$ and $81.6\%$, respectively. Ablations are detailed in Section. 4.1. We plot the training curve of the iLLaMA-T/16 in Figure 3(b), using a constant scheme with 50 cutoff epochs. When soft mask ends, we observe that although there was a sharp drop in accuracy, the model ends up achieving better performance. Similar case of the iLLaMA-B/16 are shown in Appendix E. Additionally, we discover that a lower learning rate warmup helps iLLaMA-T/16 achieve $75.0\%$ accuracy, by using a constant scheme with 50 cutoff epochs. However, such approach leads larger iLLaMA-B/16 training to fail to converge. As a result, we only used the lower learning rate warmup for iLLaMA-T/16. Notably, the final training loss within both iLLaMA-T/16 and iLLaMA-B/16 decreases when using soft masks, suggesting a mitigation of the potential underfitting concern.

### v4

@@ Abstract
This work examines whether decoder-only Transformers such as LLaMA, which were originally designed for large language models (LLMs), can be adapted to the computer vision field. We first "LLaMAfy" a standard ViT step-by-step to align with LLaMA’s architecture, and find that directly applying a causal mask to the self-attention brings an attention collapse issue, resulting in the failure to the network training. We suggest to reposition the class token behind the image tokens with a post-sequence class token technique to overcome this challenge, enabling causal self-attention to efficiently capture the entire image’s information. Additionally, we develop a soft mask strategy that gradually introduces a causal mask to the self-attention at the onset of training to facilitate the optimization behavior. The tailored model, dubbed as image LLaMA (iLLaMA), is akin to LLaMA in architecture and enables direct supervised learning. Its causal self-attention boosts computational efficiency and learns complex representation by elevating attention map ranks. iLLaMA rivals the performance with its encoder-only counterparts, achieving 75.1% ImageNet top-1 accuracy with only 5.7M parameters. Scaling the model to $\sim$310M and pre-training on ImageNet-21K further enhances the accuracy to 86.0%. Extensive experiments demonstrate iLLaMA’s reliable properties: shape-texture bias, calibration, quantization compatibility, ADE20K segmentation and CIFAR transfer learning. We hope our study can kindle fresh views to visual architectures in the wave of LLMs and inspire the development of unified multimodal models. Pre-trained models and codes are available https://github.com/techmonsterwang/iLLaMA.
@@ 1 Introduction
The year 2024 saw the meteoric rise of large language models (LLMs), as well as the 4th anniversary of the Vision Transformer (ViT) dosovitskiy2020image . Born in 2020, ViT was influenced by the prevailing encoder-only text Transformers at the time, such as BERT devlin2018bert , etc. Accordingly, ViT is allowed to borrow the encoder-only design, i.e., self-attention do not use any causal mask. As a result, advanced vision backbones touvron2021training ; wang2021pyramid ; guo2022cmt and training paradigms bao2021beit ; he2022masked have followed such convention by default.
Meanwhile, the development of text Transformers did not stand still. A series of LLMs with a decoder-only architecture (e.g., LLaMA touvron2023llama ), have sparked a new wave. Pre-trained decoder-only Transformers have demonstrated remarkable performance in diverse textual touvron2023llama2 and multimodal tasks liu2023visual ; liu2023improved ; zhu2023minigpt ; chen2023minigpt . In this context, designing unified architectures for language and vision is a promising direction. Specifically, unified models employ shared types of operators to process both visual and textual data, reducing the cost of hardware implementation. Moreover, compared to using separate models for image and text, unified models fuyu-8b ; li2023otterhd simplify the inference process by handling different modalities simultaneously, thereby improving practical efficiency.
@@
Toward this goal, we took the initial step of exploring whether decoder-only Transformers can hold its effectiveness in the unimodal vision domain. In this study, we demonstrate that through straightforward supervised learning, LLaMA architecture itself can process input images with simple yet crucial modifications. We start by modifying a standard encoder-only ViT (e.g., ViT-T/16), progressively adapting its components to align with those in LLaMA. In practice, we observe an attention collapse issue, i.e., the training loss fails to converge by directly adding a causal mask to the attention map. The causal mask restricts the class token from accessing the image’s global information, thereby hindering the optimization of the training loss. To this end, we propose a post-sequence class token technique, repositioning the class token to the end of image tokens (details in Section 3.3). As a result, causal mask can keep the attention score between the class token and others, allowing the model to optimize stably. We also evaluate the advantages of the causal self-attention in reducing computational complexity and enhancing the attention map rank.
@@
Figure 1: Left: iLLaMA architecture. Right: our design roadmap. Colored and gray bars represent the results of the tiny and base regimes, with the red line depicting the training loss of the tiny regime. iLLaMA strives to process visual tokens using standard LLaMa components, e.g., causal self-attention. The proposed PS [cls] and soft mask strategy help overcome training challenges. Block details of ViT dosovitskiy2020image , VisionLLaMA chu2024visionllama , and our iLLaMA is compared in Figure 5 in Appendix A.
@@
Equipped with these modifications, we propose a decoder-only vision Transformer with causal self-attention inside, dubbed image LLaMA (iLLaMA), as shown in Figure 1. We conduct a thorough evaluation of iLLaMA’s properties, including ImageNet-1K classification deng2009imagenet , calibration, shape-texture bias, quantization compatibility, ADE20K semantic segmentation zhou2019semantic , and CIFAR transfer learning krizhevsky2009learning . Experimental results show that iLLaMA delivers favorable and reliable performance to its encoder-only counterparts (i.e., ViT, VisionLLaMA), while maintaining a pure decoder design. More importantly, a spectral analysis on the attention map empirically shows that compared to bi-directional counterparts, causal self-attention has a higher rank (see Figure 4), which allows for learning complex image representation. The contribution of our work can be summarized as follows:
  • We explore the adaption of the LLaMA decoder to visual tasks and identify the attention collapse issue caused by causal self-attention. To address this, we introduce the PS [CLS] method.
  • We propose a soft mask strategy to optimize causal self-attention and analyzed the improvement in the rank of the causal attention map.
  • We develop a series of iLLaMA models and empirically validate its performance on ImageNet, along with practical properties such as quantization compatibility, calibration, shape-texture bias.
We hope our work to inspire a re-evaluation of vision backbone design in the era of LLMs and provide fresh insights for their architectural unification.
@@ 3 A Roadmap: Solving Attention Collapse and Optimization Improvement
This section introduces the design roadmap of iLLaMA. As we aim to adapt LLMs to vision, we choose LLaMA touvron2023llama and ViT dosovitskiy2020image as language and vision baselines due to their successful practices. The trajectory can be divided into two dimensions, i.e., architecture (Section 3.1-3.4) and training techniques (Section 3.5-3.6). First, we focus on block designs including 1) feed foward network, 2) normalization layer, 3) self-attention, 4) positional embedding, illustrated in Figure 1. Next, we study training techniques and develop a soft mask strategy to facilitate optimization. Finally, we provide an analysis in terms of efficiency and attention map rank (Section 3.7). We start with ViT-T/16 and ViT-B/16 with around 5.7M and 86.4M parameters, respectively, and gradually replace the corresponding components with those from LLaMA. We conduct experiments on ImageNet-1K deng2009imagenet , following the training recipe adopted from liu2023dropout (details in Table 12 of Appendix C.1). Considering the differences between visual perception and text generation tasks, we maintain ViT’s non-autoregressive manner in our network. Each step change and the corresponding results are reported in Table 17 of Appendix D.
@@ 3.1 Feed Forward Network (FFN)
FFN in Transformer are implemented differently in ViT and LLaMa, i.e., multi-layer perceptron (MLP) and SwiGLU shazeer2020glu . MLP consists of two sequential linear mappings, with a GELU hendrycks2016gaussian function inserted. Meanwhile, SwiGLU combines three linear mappings, integrating a SiLU hendrycks2016gaussian ; elfwing2018sigmoid ; ramachandran2017searching function. This structure allows for the modulation of high-dimensional features through a gating mechanism. We substitute the Transformer’s MLPs with SwiGLUs, while maintaining comparable computational cost. As shown in Figure 1, this improves performance from $73.8\%$ to $74.3\%$, and from $81.3\%$ to $82.0\%$ for the ViT-T/16 and ViT-B/16 regime. This highlights SwiGLU’s effectiveness not only in language models but also in vision, inspiring further exploration of other components.
@@ 3.2 Normalization Layer
Transformers need normalization layer for stable training, i.e., layer normalization (LN) ba2016layer in ViT and root mean square layer normalization (RMSNorm) zhang2019root in LLaMA, respectively. We replace all LNs with RMSNorms in our network and empirically observed that the accuracy of the ViT-T/16 regime increased from $74.3\%$ to $74.5\%$. However, similar improvements in precision were not observed in the ViT-B/16 regime (from $82.0\%$ to $81.7\%$). Nonetheless, compared to LN, RMSNorm removes the shift term computation, bringing simplicity to the network touvron2023llama2 ; vicuna2023 ; roziere2023code ; jiang2023mistral .
@@
 (a) causal mask      (b) causal mask w/ PS [cls] (ours)    (c) modified causal mask (ablation)
Figure 2: (a) mask in causal self-attention. (b) mask in causal self-attention with our post-sequence class token (PS [cls]) method. (c) modified causal mask. Their ablation results are shown in Table 1.
@@ 3.3 Causal Self-Attention Leads to Attention Collapse
Attention collapse issue. As a common practice for Transformer decoders, the key component for causal self-attention is the causal mask, i.e., a lower triangular mask matrix, illustrated in Eq. 1 and Figure 2(a). With such, each token can get the attention score of all its previous ones. We add the causal mask to our network via a non-autoregressive way. The reason is that visual perception tasks, unlike text generation, require only inference once. As a result, we observe that the training loss fails to converge in both ViT-T/16 and ViT-B/16 regimes (line 1 in Table 1). We posit that such issue stems from the influence of the lower triangular matrix, which prevents the class token from "seeing" other image tokens. As illustrated in Figure 2(a), when the class token is positioned at the start of the patch embedding, its attention score for all other image tokens gets zero due to a causal mask. We term this occurrence as the attention collapse issue, which leads to a loss of connection between the class token and other image patches, thereby hindering the optimization of the network.
@@
To test our hypothesis about the reason of the attention collapse issue, we also explore a mask setting in Figure 2(c). In this setting, we do not change the position of the class token. Instead, we unmask the first row of the mask (i.e., attention score of the class token) on the basis of the causal self-attention, termed as "modified causal mask". Ablation results (line 3 in Table 1) shows that both settings can solve the attention collapse issue as expected, and the "modified causal mask" leads to a better $72.5\%$ accuracy for ViT-T/16 regime, validating our hypothesis about the reason. Although the results do not surpass the performance of bi-directional counterpart, they demonstrate the potential for optimizing causal self-attention in a decoder-only image model. We also observe that the PS [cls] method yields higher accuracy with a slightly larger training loss for ViT-B/16 regime, suggesting lower overfitting.
@@ 3.4 Positional Embedding
A standard ViT use learnable positional embedding (LPE) to preserve positional information, typically adding it directly to the patch embedding. Meanwhile, rotary positional embedding (RoPE) su2024roformer is widely employed in LLMs touvron2023llama ; touvron2023llama2 , which functions in the attention of each block. We first use RoPE alone, which boosts the accuracy of ViT-T/16 and ViT-B/16 regimes to $72.6\%$ and $81.2\%$, from $71.9\%$ and $80.6\%$, respectively. The encouraging results illustrate that the concepts of "position" in image and text do not exist an inherent gap. Since LPE functions only once before all Transformer blocks, keeping it does not disrupt the alignment with LLaMA within each block. Thus, we reintroduce the LPE, which improves the accuracy of ViT-T/16 regime from $72.6\%$ to $73.2\%$, suggesting that the two positional embeddings are not redundant but rather synergistic, enhancing network performance.
@@ 3.5 Data Augmentation
Mixup zhang2017mixup and cutmix yun2019cutmix that we used to train our iLLaMA (0.8 and 1.0), are borrowed from DeiT touvron2021training ’s recipe. Unlike the bi-directional self-attention used in DeiT, causal self-attention affects the connection between image tokens. Meanwhile, these two hyper-parameters affect the content of the input image, which further influences the subsequent embedding. Thus, we reevaluate their impact on iLLaMA optimization. Specifically, we discover that a combination of 0.1 mixup and 0.1 cutmix improves the performance of the iLLaMA-T/16 from $73.2\%$ to $74.3\%$, whereas a combination of 0.95 and 1.0 leads the iLLaMA-B/16 to a $81.3\%$ accuracy. Other ablations are detailed in Section 4.1.
@@ 3.6 Soft Mask Strategy: Optimization Improvement
Figure 3: (a) Soft mask gradually transitions from a bi-directional mask into a causal mask during training through a constant or linear schedule. (b) Ablation results of training loss and test accuracy.
where $i,j\in[1,N]$, ${\bf S}\in\mathbb{R}^{N\times N}$ denotes the soft mask, which is defined as a linear combination of a bi-directional mask ${\bf B}$ and a causal mask ${\bf C}$. $\alpha$ is the hyper-parameter controlling the mask configuration, i.e., soft mask degenerates into ${\bf B}$ or ${\bf C}$ when $\alpha=1$ or $\alpha=0$, respectively. As illustrated in Figure 3(a), $\alpha$ involves three related hyper-parameters: 1) scheme: how $\alpha$ drops from 1 to 0: we try a linear or a constant scheme. 2) cutoff epochs: when will $\alpha$ drops to 0. 3) learning rate (lr) warmup he2016deep ; goyal2017accurate : this hyper-parameter overlaps with the duration of soft mask. We initially set the lr warmup epochs at 50, consistent with previous settings. When using a linear scheme with 50 and 25 cutoff epochs, we observe an improvement in performance for both iLLaMA-T/16 and iLLaMA-B/16 models, reaching $74.9\%$ and $81.6\%$ from $74.3\%$ and $81.3\%$, respectively. Ablations results are detailed in Section. 4.1. To intuitively observe the impact of soft mask, we plot the training curve of the iLLaMA-T/16 in Figure 3(b), using a constant scheme with 50 cutoff epochs. When soft mask ends, we observe that although there was a sharp drop in accuracy, the model ends up achieving better performance. Similar case of the iLLaMA-B/16 are shown in Figure 6 of Appendix F. Additionally, we discover that a lower learning rate warmup helps iLLaMA-T/16 achieve $75.0\%$ top-1 accuracy, by using a constant scheme with 50 cutoff epochs. Therefore, we use this warmup method for iLLaMA-T/16. Notably, the final training loss within both iLLaMA-T/16 and iLLaMA-B/16 decreases when using soft masks, suggesting a mitigation of potential underfitting concerns.